# Preconditioned Score-based Generative Models

### v2

@@ 5.2 Constructing image preconditioners
In this section, we design the frequency $M_{f}$ and pixel $M_{p}$ preconditioners. We denote an image training dataset $\mathcal{D}=\{\mathbf{x}^{(i)}\in\mathbb{R}^{C\times H\times W},i=1,\dots,N\}$, where $C,H$ and $W$ are the channel number, height and width of images, respectively.
@@ 5.2.1 Frequency preconditioner $M_{f}$
where $\alpha\in[1,+\infty)$ is the normalization parameter. The intuition is that, the mask $M_{f}$ shrinks the frequency coordinates with larger scales more than the others, so that the scales of all coordinates can become similar. This mitigates the ill-conditioned issue in the frequency space as mentioned earlier.
@@ 5.2.2 Pixel preconditioner $M_{p}$
$M_{p}[\mathbf{x}]=\mathbf{x}\bullet R_{p},\quad\forall\mathbf{x}\in\mathbb{R}^{C\times W\times H}$ (35)
@@
with the normalization parameter $\alpha$ same as in Eq. (34).
For the tasks without clear pixel structure characteristics, we do not apply this pixel preconditioner by setting all the elements of $R_{p}$ to $1$ (i.e., an identity operator).
@@ 5.2.3 Computational complexity
Only a small set of images are needed to estimate the statistics for Eq. (33) and Eq. (36). In practice, we use $200$ random images with marginal cost. For the computational complexity of PDS during sampling, the major overhead is from FFT [28] and its inverse with the complexity of $O(CHW(\log H+\log W))=O(d(\log H+\log W))$. This is neglectable w.r.t the original cost (see Table V).
@@ 6.1 Gradient descent perspective
We can understand the preconditioning on the Langevin dynamics from the gradient descent perspective. Given the target data distribution $p_{\mathbf{x}^{\ast}}$, we define the energy function (a.k.a. surprise in cognitive science [55])
$E(\mathbf{x})=-\log p_{\mathbf{x}^{\ast}}(\mathbf{x}),\quad\forall\mathbf{x}\in\mathbb{R}^{d}$ (38)
Substituting the definition of $E$ into Eq. (10), we have
$\mathbf{x}_{t}=\mathbf{x}_{t-1}-\frac{\epsilon_{t}^{2}}{2}\bigtriangledown_{\mathbf{x}}E(\mathbf{x}_{t-1})+\epsilon_{t}\mathbf{z}_{t},1\leq t\leq T.$ (39)
Clearly, the Langevin dynamics is implementing the gradient descent algorithm for minimizing the energy of the sample point $\mathbf{x}_{t}$, along with an additive noise term $\mathbf{z}_{t}$. This optimizer is called stochastic gradient Langevin dynamics (SGLD) [24]. Unlike stochastic gradient descent, the randomness of SGLD comes from the noise $\mathbf{z}_{t}$, instead of drawing from the training set. Besides, SGMs sequentially reduce the energy ($-\log p_{\mathbf{x}^{\ast}}(\mathbf{x})$) of a sample $\mathbf{x}_{t}$ via SGLD.
@@ 6.2 Skew-symmetric operator
In the gradient descent perspective, $S\bigtriangledown_{\mathbf{x}}E(\mathbf{x}_{t-1})dt$ of Eq. (24) is known as solenoidal term, since it corresponds to the direction along which the energy does not change (i.e., $\langle\bigtriangledown_{\mathbf{x}}E(\mathbf{x}_{t-1}),S\bigtriangledown_{\mathbf{x}}E(\mathbf{x}_{t-1})\rangle=0$). Seemingly it imposes no effect on acceleration. But several studies [56, 57, 58, 59] suggest that it can help reach the steady state distribution faster.
For completeness, we disucss the possibility of the skew-symmetric operator $S$ as an accelerator for the Langevin dynamics. According to [56], under the regularity condition that $\left|\mathbf{x}_{t}\right|$ does not reach the infinity in a finite time, the convergence of an autonomous (the right side of the equation does not contain time explicitly) Langevin dynamics
$d\mathbf{x}_{t}=\frac{\epsilon_{t}^{2}}{2}\bigtriangledown_{\mathbf{x}_{t}}\log p_{\mathbf{x}^{\ast}}(\mathbf{x}_{t})dt+\epsilon_{t}d\mathbf{w}_{t},$ (40)
@@
$d\mathbf{x}_{t}=\frac{\epsilon_{t}^{2}}{2}\bigtriangledown_{\mathbf{x}_{t}}\log p_{\mathbf{x}^{\ast}}(\mathbf{x}_{t})dt+C(\mathbf{x}_{t})dt+\epsilon_{t}d\mathbf{w}_{t},$ (41)
@@
$\bigtriangledown_{\mathbf{x}}\cdot(\frac{C(\mathbf{x})}{p_{\mathbf{x}^{\ast}}(\mathbf{x})})=0.$ (42)
@@ 6.3 Parameter selection
We investigate how to set properly $\alpha$ of Eq. (37) and Eq. (34). Increasing $\alpha$ will weaken the preconditioning effect; When $\alpha\rightarrow+\infty$, $M_{f}$ and $M_{p}$ both become the identity operator. Empirically, we find that for higher iterations $T$, higher $\alpha$ is preferred (Table I). We aim to quantify the relation between $\alpha$ and $T$.
By Thm. 3, the preconditioned reverse diffusion (Eq. (5.1)) reverses the transformed forward process (Eq. (26)), instead of the original one (Eq. (19)) used in model training. The deviation between Eq. (26) and Eq. (19) brings challenge for the model to fit the score function, as the noisy input distribution deviates from what is exposed to model training at each step. We assume the error per iteration grows linearly with this deviation, proportional to
$\left\|M_{p}[M_{f}[\cdot]]-I[\cdot]\right\|,$ (43)
@@
With Eq. (34) and (37), in case of large
@@
$T\sqrt{d}[(1+O(\frac{1}{\alpha}))(1+O(\frac{1}{\alpha}))-1].$ (45)
Omitting high order terms, it could be simplified as
$T\sqrt{d}[(1+O(\frac{1}{\alpha}))(1+O(\frac{1}{\alpha}))-1]=O(1),$ (46)
@@
$(1+\frac{c_{1}}{\alpha})(1+\frac{c_{2}}{\alpha})-1=a\frac{1}{T}+b,$ (47)
@@ 7 Experiments
In this section, we first show that PDS can accelerate the sampling process of the state-of-the-art SGMs while maintaining the sampling quality and provide parameter analysis. We use CIFAR-10 [60], CelebA [61] at a resolution of $64\times 64$, LSUN (bedroom and church) [62] at a resolution of $256\times 256$, and FFHQ [21] at a resolution of $1024\times 1024$. Note that for all these datasets, the image height and width are identical, i.e., $H=W$. The CelebA and FFHQ datasets are human facial images, with strong pixel-wise structural characteristics. For them, we hence apply both frequency and pixel filter for these two datasets. For other datasets, we only apply the frequency preconditioner. We adopt the same preprocessing as [4]. We evaluate PDS on the state-of-art SGM i.e., NCSN++ [4]. For fair comparison, we only replace the vanilla sampling method (Alg. 1) with our PDS (Alg. 2) whilst keeping the remaining unchanged during inference for a variety of sampling iterations, without model retraining and other alteration. Whenever reducing the iterations from the original one $T$ to $T/c$ with a factor $c$ we simultaneously expand the step size from the original one $\epsilon_{t}$ to $\sqrt{c}\epsilon_{t}$ for $t=1,\ldots,T$ for consistent accumulative update of the score function and consistent variance of the added noises. We use the public codebase of [4]. We use the released checkpoints for all SGMs except NCSN++ on CelebA ($64\times 64$), for which no released checkpoint, and we train by ourselves with the released codes instead.
@@ 7.1 Evaluating on sampling quality
We show that PDS can accelerate the sampling process of SGMs (NCSN++) for generating samples on datasets across various resolutions. See Table I the setting of parameter $\alpha$ for different iterations $T$ by our linear relation (Sec. 6.3).
TABLE I: Parameters $\alpha$ of PDS used for constructing preconditioner on NCSN++ [4] under different iterations $T$.
@@
CIFAR-10 and CelebA. In Table II and Table III, we evaluate the PDS with NCSN++ on CIFAR-10 ($32\times 32$) and CelebA ($64\times 64$) under various iterations $T$. We use FID [63] score as metric. We compare PDS with representative DDPMs (DDPM [35], DDIM [30], Analytic-DDPM [31], Analytic-DDIM [31], and PNDM [38]). We provide visual samples on CIFAR-10 and CelebA in Appendix.
Although achieving on-par performance with DDPMs given large iterations, NCSN++ degrades more significantly as the iteration decreases. This can be solved by our PDS whilst further improving large-iteration performance and achieving the best results in most low-iteratation cases. As shown in Table IV, our PDS achieves a FID score of 1.99 on CIFAR-10, surpassing prior GANs, DDPMs and SGMs.
TABLE II: FID scores under different iterations $T$ on CIFAR-10.
Iterations $T$ NCSN++ [4] NCSN++ w/ PDS DDPM [35] DDIM [30] Analytic-DDPM [31] Analytic-DDIM [31] PNDM [38]
50 306.91 4.90 66.28 7.74 7.25 4.04 3.95
100 29.39 2.90 31.36 6.08 5.40 3.55 3.72
200 4.35 2.53 12.96 5.07 4.01 3.39 3.48
400 2.40 2.19 4.86 4.61 3.62 3.50 3.57
1000 2.23 1.99 3.17 4.13 4.03 3.74 3.70
TABLE III: FID scores under different iterations $T$ on CelebA ($64\times 64$).
Iterations $T$ NCSN++ [4] NCSN++ w/ PDS DDPM [35] DDIM [30] Analytic-DDPM [31] Analytic-DDIM [31] PNDM [38]
50 438.84 13.48 115.69 9.33 11.23 6.13 5.69
100 433.98 3.49 25.65 6.60 8.08 4.29 4.03
200 292.51 3.21 9.72 4.96 6.51 3.46 3.04
400 7.09 2.95 3.95 4.15 5.87 3.38 3.07
1000 3.33 3.25 3.16 3.40 5.21 3.13 2.99
TABLE IV: FID comparison on CIFAR-10.
@@
GAN BigGAN [64] 14.73
StyleGAN2-ADA [65] 2.42
DDPM DDPM [35] 3.17
DDIM [30] 4.13
Analytic-DDPM [31] 3.63
Analytic-DDIM [31] 3.39
PNDM [38] 3.26
DDPM++ [4] 2.91
SGM NCSN++ [4] 2.23
CLD-SGM [29] 2.23
LSGM [34] 2.10
INDM [66] 2.28
@@
LSUN [62]. We test PDS with NCSN++ to generate bedroom and church images at a resolution of $256\times 256$. As shown in Fig. 3, our PDS can prevent the output images from being ruined by heavy noises under low iterations. See more visualization in Appendix.
Figure 3: Sampling on LSUN (church and bedroom) ($256\times 256$). Top: NCSN++ [4] Bottom: NCSN++ with PDS. Both use $80$ iterations and $\alpha=2.2$. More samples in Appendix.
FFHQ [21]. We further test PDS to generate high-resolution facial images at a resolution of $1024\times 1024$. Similarly under acceleration the original sampling method suffers from heavy noises and fails to produce recognizable human faces. Our PDS maintains the sampling quality using as few as 66 iterations (Fig. 1 and Fig. 10). ImageNet [67] and COCO [68]. We further investigate the effect of PDS on NCSN++ in generating samples on ImageNet and COCO respectively. As the first ImageNet and COCO evaluation for VE-SDE models, we train a NCSN++ [4] by ourselves at the resolution of 64$\times$64 as the baseline. We apply the same hyper-parameters of optimizer and learning schedule as [4]. We train the model for 110k iterations for ImageNet and 600k for COCO. The batch size is set at 256 for both datasets. As shown in Fig. 4, our PDS yields consistently acceleration effect across a variety of sampling iterations. We show image samples in Fig. 7 and Fig. 8 of Appendix.
Figure 4: The FID score of NCSN++ [4] with and without PDS on ImageNet and COCO under various iterations $T$.

### v3

@@ 5.3 Constructing image preconditioners
In this section, we design the frequency $M_{f}$ and pixel $M_{p}$ preconditioners. We denote an image training dataset as $\mathcal{D}=\{\mathbf{x}^{(i)}\in\mathbb{R}^{C\times H\times W},i=1,\dots,N\}$, where $C,H$ and $W$ are the channel number, height and width of images, respectively.
@@ 5.3.1 Frequency preconditioner $M_{f}$
where $\alpha\in[1,+\infty)$ is the normalization parameter. The intuition is that the mask $M_{f}$ shrinks the frequency coordinates with larger scales more than the others so that the scales of all coordinates can become similar. This mitigates the ill-conditioned issue in the frequency space, as mentioned earlier.
@@ 5.3.2 Pixel preconditioner $M_{p}$
$M_{p}[\mathbf{x}]=\mathbf{x}\bullet R_{p},\quad\forall\mathbf{x}\in\mathbb{R}^{C\times W\times H}$
@@
with the normalization parameter $\alpha$ same as in Eq. (29). For the tasks without clear pixel structure characteristics, we do not apply this pixel preconditioner by setting all the elements of $R_{p}$ to $1$ (i.e., an identity operator).
@@ 5.3.3 Computational complexity
Only a small set of images is needed to estimate the statistics for Eq. (28) and Eq. (30). In practice, we use $200$ random images with marginal cost. For the computational complexity of PDS during sampling, the major overhead is from FFT (Brigham, 1988) and its inverse with the complexity of $O(CHW(\log H+\log W))=O(d(\log H+\log W))$. This is neglectable w.r.t the original cost (see Table 8).
@@ 6.1 Gradient descent perspective
We can understand the preconditioning of the Langevin dynamics from the gradient descent perspective. Given the target data distribution $p_{\mathbf{x}^{\ast}}$, we define the energy function (a.k.a. surprise in cognitive science (Friston, 2010)) as $E(\mathbf{x})=-\log p_{\mathbf{x}^{\ast}}(\mathbf{x}),\quad\forall\mathbf{x}\in\mathbb{R}^{d}$. Substituting the definition of $E$ into Eq. (8), we have
$\mathbf{x}_{t}=\mathbf{x}_{t-1}-\frac{\epsilon_{t}^{2}}{2}\bigtriangledown_{\mathbf{x}}E(\mathbf{x}_{t-1})+\epsilon_{t}\mathbf{z}_{t},1\leq t\leq T.$
Clearly, the Langevin dynamics is implementing the gradient descent algorithm for minimizing the energy of the sample point $\mathbf{x}_{t}$, along with an additive noise term $\mathbf{z}_{t}$. This optimizer is called stochastic gradient Langevin dynamics (SGLD) (Welling and Teh, 2011). Unlike stochastic gradient descent, the randomness of SGLD comes from the noise $\mathbf{z}_{t}$, instead of drawing from the training set. Besides, SGMs sequentially reduce the energy ($-\log p_{\mathbf{x}^{\ast}}(\mathbf{x})$) of a sample $\mathbf{x}_{t}$ via SGLD.
@@ 6.2 Skew-symmetric operator
In the gradient descent perspective, $S\bigtriangledown_{\mathbf{x}}E(\mathbf{x}_{t-1})dt$ of Eq. (20) is known as the solenoidal term, since it corresponds to the direction along which the energy does not change (i.e., $\langle\bigtriangledown_{\mathbf{x}}E(\mathbf{x}_{t-1}),S\bigtriangledown_{\mathbf{x}}E(\mathbf{x}_{t-1})\rangle=0$). Seemingly it imposes no effect on acceleration. Several studies (Hwang et al., 2005; Ottobre, 2016; Rey-Bellet and Spiliopoulos, 2015; Lelievre et al., 2013) suggest that it can help reach the steady state distribution faster.
For completeness, we discuss the possibility of using the skew-symmetric operator $S$ for accelerating the Langevin dynamics. According to (Hwang et al., 2005), under the regularity condition that $\left|\mathbf{x}_{t}\right|$ does not reach the infinity in a finite time, the convergence of an autonomous (the right side of the equation does not contain time explicitly) Langevin dynamics
$d\mathbf{x}_{t}=\frac{\epsilon_{t}^{2}}{2}\bigtriangledown_{\mathbf{x}_{t}}\log p_{\mathbf{x}^{\ast}}(\mathbf{x}_{t})dt+\epsilon_{t}d\mathbf{w}_{t},$
@@
$d\mathbf{x}_{t}=\frac{\epsilon_{t}^{2}}{2}\bigtriangledown_{\mathbf{x}_{t}}\log p_{\mathbf{x}^{\ast}}(\mathbf{x}_{t})dt+C(\mathbf{x}_{t})dt+\epsilon_{t}d\mathbf{w}_{t},$
@@
$\bigtriangledown_{\mathbf{x}}\cdot(\frac{C(\mathbf{x})}{p_{\mathbf{x}^{\ast}}(\mathbf{x})})=0.$
@@ 6.3 Parameter selection
We investigate how to set properly $\alpha$ of Eq. (31) and Eq. (29). Increasing $\alpha$ will weaken the preconditioning effect; When $\alpha\rightarrow+\infty$, $M_{f}$ and $M_{p}$ both become the identity operator. Empirically, we find that for higher iterations $T$, higher $\alpha$ is preferred (Table 1). We aim to quantify the relation between $\alpha$ and $T$. By Thm. 4.2, the preconditioned reverse diffusion (Eq. (26)) reverses the transformed forward process (Eq. (22)), instead of the original one (Eq. (16)) used in model training. The deviation between Eq. (22) and Eq. (16) brings a challenge for the model to fit the score function, as the noisy input distribution deviates from what is exposed to model training at each step. We assume the error per iteration grows linearly with this deviation, proportional to
$\left\|M_{p}[M_{f}[\cdot]]-I[\cdot]\right\|,$
@@
With Eq. (29) and (31), in case of large $\alpha$ we have
@@
$T\sqrt{d}[(1+O(\frac{1}{\alpha}))(1+O(\frac{1}{\alpha}))-1].$
Omit high-order terms, it could be simplified as
$T\sqrt{d}[(1+O(\frac{1}{\alpha}))(1+O(\frac{1}{\alpha}))-1]=O(1).$ (33)
@@
$(1+\frac{c_{1}}{\alpha})(1+\frac{c_{2}}{\alpha})-1=a\frac{1}{T}+b.$ (34)
@@ 7 Experiments
In this section, we first show that PDS can accelerate the sampling process of state-of-the-art SGMs while maintaining the sampling quality and providing parameter analysis. We use CIFAR-10 (Krizhevsky and Hinton, 2009), CelebA (Liu et al., 2015) at a resolution of $64\times 64$, LSUN (bedroom and church) (Yu et al., 2015) at a resolution of $256\times 256$, and FFHQ (Karras et al., 2019) at a resolution of $1024\times 1024$. Note that for all these datasets, the image height and width are identical, i.e., $H=W$. The CelebA and FFHQ datasets are human facial images, with strong pixel-wise structural characteristics. Hence, we apply both frequency and pixel filters to these two datasets. For other datasets, we only apply the frequency preconditioner. We adopt the same preprocessing as (Song et al., 2021b). We evaluate PDS on the state-of-art SGM i.e., NCSN++ (Song et al., 2021b). For a fair comparison, we only replace the vanilla sampling method (Alg. 1) with our PDS (Alg. 2) whilst keeping the remaining unchanged during inference for a variety of sampling iterations, without model retraining and other alteration. Whenever reducing the iterations from the original one $T$ to $T/c$ with a factor $c$ we simultaneously expand the step size from the original one $\epsilon_{t}$ to $\sqrt{c}\epsilon_{t}$ for $t=1,\ldots,T$ for a consistent accumulative update of the score function and consistent variance of the added noises. We use the public codebase of (Song et al., 2021b). We use the released checkpoints for all SGMs except NCSN++ on CelebA ($64\times 64$), for which there is no released checkpoint, and we train by ourselves with the released codes instead. See Appendix for the details of implementation.
@@ 7.1 Evaluating on sampling quality
We show that PDS can accelerate the sampling process of SGMs (NCSN++) for generating samples on datasets across various resolutions. See Table 1 the setting of parameter $\alpha$ for different iterations $T$ by our linear relation (Sec. 6.3).
Table 1: Parameters $\alpha$ of PDS used for constructing preconditioner on NCSN++ (Song et al., 2021b) under different iterations $T$.
@@
CIFAR-10 and CelebA. In Table 2 and Table 3, we evaluate the PDS with NCSN++ on CIFAR-10 ($32\times 32$) and CelebA ($64\times 64$) under various iterations $T$. We provide visual samples on CIFAR-10 and CelebA in Appendix. The noise schedule defines how noise is progressively added to the data during the forward process. Two commonly used noise schedules are linear and cosine. Unless specified otherwise, the methods presented in Table 2, Table 3, and Table 7 employed a linear noise schedule. Since our baseline, NCSN++ (Song et al., 2021b), employs a linear schedule, we also adopt a linear schedule for the PDS evaluation.
Although achieving on-par performance with DDPMs given high iterations, NCSN++ degrades more significantly as the number of iterations decreases. Specifically, NCSN++ performs the worst among all methods at both 50 and 100 iterations. Our PDS addresses this limitation while further enhancing performance of the NCSN++ at high iteration. As demonstrated in Table 2 and Table 3, our PDS achieves the best FID scores at 100 iterations on both the CIFAR-10 and CelebA datasets, while also delivering competitive results at 50 iterations. The reason PNDM (Liu et al., 2021) outperforms our PDS at the 50 iterations is that our base method NCSN++ suffers from significant numerical instability under low-iteration conditions, which cannot be entirely resolved by our PDS. Furthermore, as shown in Table 7, our PDS achieves an FID score of 1.99 on CIFAR-10, surpassing prior competitors.
Table 2: FID scores under different iterations $T$ on CIFAR-10.
Iterations $T$ 50 100 200 400 1000
NCSN++ (Song et al., 2021b) 306.91 29.39 4.35 2.40 2.23
NCSN++ w/ PDS 4.90 2.90 2.53 2.19 1.99
DDPM (Ho et al., 2020) 66.28 31.36 12.96 4.86 3.17
DDIM (Song et al., 2020) 7.74 6.08 5.07 4.61 4.13
Analytic-DDPM (linear(Bao et al., 2022b) 7.25 5.40 4.01 3.62 4.03
Analytic-DDPM (cosine(Bao et al., 2022b) 5.50 4.45 4.04 3.96 4.31
Analytic-DDIM (linear(Bao et al., 2022b) 4.04 3.55 3.39 3.50 3.74
Analytic-DDIM (cosine(Bao et al., 2022b) 6.02 4.88 4.92 5.00 4.66
PNDM (linear(Liu et al., 2021) 3.95 3.72 3.48 3.57 3.70
PNDM (cosine(Liu et al., 2021) 3.68 3.53 3.50 3.30 3.26
Table 3: FID scores under different iterations $T$ on CelebA ($64\times 64$).
Iterations $T$ 50 100 200 400 1000
NCSN++ (Song et al., 2021b) 438.84 433.98 292.51 7.09 3.33
NCSN++ w/ PDS 13.48 3.49 3.21 2.95 3.25
DDPM (Ho et al., 2020) 115.69 25.65 9.72 3.95 3.16
DDIM (Song et al., 2020) 9.33 6.60 4.96 4.15 3.40
Analytic-DDPM (Bao et al., 2022b) 11.23 8.08 6.51 5.87 5.21
Analytic-DDIM (Bao et al., 2022b) 6.13 4.29 3.46 3.38 3.13
PNDM (Liu et al., 2021) 5.69 4.03 3.04 3.07 2.99
Table 4: Comparison of Inception scores between NCSN++ and NCSN++ with our PDS across varying iterations $T$ on CIFAR-10 and CelebA ($64\times 64$).
Dataset Methods 50 100 200 400 1000
CIFAR-10 NCSN++ (Song et al., 2021b) 1.54 6.99 9.29 9.53 9.73
NCSN++ w/ PDS 9.53 9.36 9.42 9.56 9.81
CelebA NCSN++ 1.14 1.15 2.43 2.54 2.23
NCSN++ w/ PDS 2.54 2.47 2.41 2.28 2.22
Table 5: FID score under various iterations $T$ on ImageNet ($64\times 64$).
Iterations $T$ 100 200 400 1000
NCSN++ (Song et al., 2021b) 422.54 134.17 33.48 18.93
NCSN++ w/ PDS 19.1 17.08 16.15 15.63
DDPM (Ho et al., 2020) 45.04 28.39 21.38 17.58
DDIM (Song et al., 2020) 18.09 17.84 17.74 17.73
Analytic-DDPM (Bao et al., 2022b) 18.80 17.16 16.40 16.14
Analytic-DDIM (Bao et al., 2022b) 17.73 17.49 17.44 17.57
Table 6: FID score under various iterations $T$ on COCO.
Iterations $T$ 2000 1000 500 400 200 100
NCSN++ 12.41 15.82 37.71 65.39 339.59 464.48
NCSN++ w/ PDS 11.76 12.77 13.51 13.63 15.94 19.57
Table 7: Best FID comparison on CIFAR-10.
@@
GAN BigGAN (Brock et al., 2018) 14.73
StyleGAN2-ADA (Kang et al., 2021) 2.42
DDPM DDPM (Ho et al., 2020) 3.17
DDIM Song et al. (2020) 4.13
Analytic-DDPM (linear(Bao et al., 2022b) 3.63
Analytic-DDPM (cosine(Bao et al., 2022b) 4.04
Analytic-DDIM (linear(Bao et al., 2022b) 3.39
Analytic-DDIM (cosine(Bao et al., 2022b) 4.66
PNDM (linear(Liu et al., 2021) 3.48
PNDM (cosine(Liu et al., 2021) 3.26
DDPM++ (Song et al., 2021b) 2.91
SGM NCSN++ (Song et al., 2021b) 2.23
CLD-SGM (Dockhorn et al., 2022) 2.23
LSGM (Vahdat et al., 2021) 2.10
INDM (Kim et al., 2022) 2.28
@@
LSUN (Yu et al., 2015). We test PDS with NCSN++ to generate bedroom and church images at a resolution of $256\times 256$. As shown in Fig. 4, our PDS can prevent outputs from being ruined by heavy noises under low iterations. See more visualization in the Appendix.
FFHQ (Karras et al., 2019). We test PDS to generate facial images at a resolution of $1024\times 1024$. Similarly, under acceleration the original sampling method suffers from heavy noises and fails to produce recognizable human faces. Our PDS maintains the sampling quality with 66 iterations (Fig. 1 and Fig. 21).
ImageNet (Deng et al., 2009) and COCO (Lin et al., 2014). We further investigate the effect of PDS on NCSN++ in generating image samples on ImageNet and COCO respectively. As the first ImageNet and COCO evaluation for VE-SDE models, we train an NCSN++ by ourselves at the resolution of 64$\times$64 as the baseline. As shown in Table 5, our PDS yields consistently acceleration effect across a variety of sampling iterations and produces comparable performance compared with NCSN++, DDPM (Ho et al., 2020), DDIM (Song et al., 2020), Analytic-DDPM (Bao et al., 2022b) and Analytic-DDIM (Bao et al., 2022b). Note that our PDS reaches the best FID score (15.63) among all methods. Since there is no available codebase of DDPM, DDIM, Analytic-DDPM, and Analytic-DDIM for COCO dataset generation, we compare our PDS with the baseline NCSN++. As shown in Tab. 6, our PDS also enhances the generation quality of NCSN++ across different sampling iterations on COCO. We showcase more samples in Fig. 19-22 of the Appendix.
Figure 4: Sampling on LSUN (church and bedroom) ($256\times 256$). Top: NCSN++ (Song et al., 2021b) Bottom: NCSN++ with PDS. Both use $80$ iterations and $\alpha=2.2$. More samples in Appendix.